# ActAlign: Zero-Shot Fine-Grained Video Classification via Language-Guided Sequence Alignment

### v2

@@ 4.2 Experimental Result
We begin with a random choice baseline and a SigLIP zero-shot baseline using mean-pooled frame embeddings. Following ViFi-CLIP (Rasheed et al. 2023), each video $V_{i}$ is encoded as $\bar{\mathbf{z}}_{i}=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\mathbf{z}_{i}^{t}$, where mean-pooled embeddings $\bar{\mathbf{z}}$ is compared via cosine similarity to each name embedding $\phi_{t}(c_{j})$.
We further compare against open-source video–language models, including entries from the ActionAtlas leaderboard and fine-tuned CLIP variants. Despite using no video–text supervision, our method outperforms the SigLIP baseline by +7% Top-1 and +11% Top-2 accuracy, surpassing all baselines and billion-parameter models with $\sim$8$\times$ fewer parameters. These larger models are often optimized for interactive tasks through extensive instruction tuning, rather than for classifying fine-grained actions. In contrast, image–language models are trained on large-scale image–text pairs, making them particularly well-suited for open-set recognition when paired with temporal alignment mechanisms.
@@
Configuration Top-1 (%) $\uparrow$ Top-2 (%) $\uparrow$ Top-3 (%) $\uparrow$ SigLIP (Zhai et al. 2023) (mean-pool) 22.9422.9422.94 42.2042.2042.20 63.7063.7063.70 + DTW Alignment 25.7225.7225.72 45.9945.9945.99 66.2666.2666.26 + Context Augmentation 30.0730.0730.07 52.6752.6752.67 70.4970.4970.49 + Signal Smoothing 30.2930.2930.29 53.0153.0153.01 70.2770.2770.27
@@
Prompt Description Top-1 (%) $\uparrow$ Short‐fixed 2-word, fixed 10 sub-actions 27.0627.0627.06 Context‐rich context-rich, variable-length 30.29
@@
Table 4 shows LLM-generated subaction sequences for two Figure skating tactics under our context-rich prompting. The scripts highlight ordered, salient steps enabling precise temporal matching.
Tactic-Subaction Script Biellmann Spin 1. Begins upright spin on one foot with arms extended and free leg behind, 2. Gradually pulls free leg upward behind the back using both hands, 3. Raises the free leg above head level while arching the back dramatically, 4. Grasps the blade of the free skate with both hands overhead, 5. Extends spinning leg vertically while maintaining centered spin on skating foot, 6. Maintains high-speed rotation with body in extreme vertical split position Flying Camel Spin 1. Skater glides forward with arms extended and knees bent in preparation, 2. Performs a powerful jump off the toe pick while swinging free leg upward, 3. Rotates mid-air with body extended horizontally like a ’T’ shape, 4. Lands on one foot directly into a camel spin position with torso parallel to ice, 5. Extends free leg backward and arms outward while spinning on the skating leg, 6. Maintains fast, centered rotation in the horizontal camel position
Table 4: LLM-generated sub-action scripts for figure skating tactics. Shown for the Biellmann Spin and Flying Camel Spin examples in Figure 4, these sequences are generated using context-rich prompting ($T=0.2$) and provide semantically detailed, temporally ordered steps for alignment in our zero-shot framework.
@@ 5 Conclusion
We show that contrastive image–language models establish a surprisingly strong baseline for zero-shot fine-grained video classification, even when used with simple mean-pooling. To fully leverage this capability in extremely fine-grained settings, we propose ActAlign, a novel zero-shot, domain-general, and model-agnostic framework that revisits the classic Dynamic Time Warping (DTW) algorithm to cast video classification as a sequence-alignment problem. By aligning video frames with LLM-generated sub-action scripts, ActAlign introduces temporal structure into contrastive models without requiring any video–text training or fine-tuning. Evaluated on the highly challenging and diverse ActionAtlas (Salehi et al. 2024) benchmark, our method achieves state-of-the-art performance, outperforming both CLIP-style baselines and billion-parameter video–language models.
@@ Appendix B Implementation Details and Inference Pipeline
Model Settings
We use the publicly available SigLIP model from Google (siglip-so400m-patch14-384), which supports variable input resolutions. Images are embedded using a ViT-based vision encoder with a patch size of $\times 14$ 14 and an output embedding dimension of $d=384$. Sub-actions are encoded using the corresponding frozen SigLIP text encoder. All embeddings are $\ell_{2}$-normalized before computing cosine similarities.
Runtime Performance
@@
Inference Pipeline Overview.
@@ Appendix C More Related Works
Image–Language Transfer for Video Tasks. Image–language models pretrained on large-scale image–text pairs exhibit strong open-set recognition capabilities. This motivates their adaptation for video tasks. Common strategies include adding temporal modules (Wang, Xing, and Liu 2022; Ni et al. 2022; Rasheed et al. 2023; Liu, Li, and Yu 2024) or introducing learnable prompts (Ju et al. 2022; Lin et al. 2022). However, as noted by (Rasheed et al. 2023), fine-tuning often compromises the model’s open-set generalization. To preserve zero-shot capabilities, recent work explores training-free adaptations using temporal prompting methods (Ahmad, Chanda, and Rawat 2024; Phan et al. 2024; Yu et al. 2025). For example, TEAR (Bosetti et al. 2025) compares mean-pooled frame embeddings with LLM-generated action descriptions for each class. While effective, these approaches either degrade generalization through fine-tuning or ignore temporal structure via mean-pooling - limiting performance on fine-grained and zero-shot video classification.
@@ Appendix D Prompt Templates and Generation Settings
This setting produces terse outputs (e.g., “tie arms,” “deliver knee”) that often lack domain-specific context. The descriptions are directly adopted from the orginal ActionAtlas (Salehi et al. 2024) dataset.
Template
@@
This setting encourages richer descriptions with more related keywords grounded in the activity domain. Context-rich prompting strategy significantly improves textual grounding during DTW alignmnet.
Template
@@ D.3 Examples
Table 5 compares the two prompting strategies for an skateboarding trick example. The corresponding affinity matrices are shown in Figure 6.
@@ Appendix E Dataset Statistics
Table 6 summarizes key statistics of the ActionAtlas (Salehi et al. 2024) dataset, which we extended to enable alignmnet-based classification. The dataset contains 898 video clips spanning 558 unique tactics across 56 sports domains. Each video is paired with 4 to 6 fine-grained candidate actions (5.25 in average per video). Videos vary significantly in length, with a mean of 173 frames (30 FPS), ranging from extremely short clips (3 frames) to multi-second sequences (up to 900 frames). Notably, ActionAtlas is the most diverse and challenging fine-grained action recognition benchmark to the best of our knowledge. It poses a major challenge in fine-grained recognition: human top-1 accuracy is just 61.64%.
@@
Figure 8: Our pipeline uses an LLM to generate structured sub-action sequences for each fine-grained candidate action in ActionAtlas (Salehi et al. 2024) data sample. This structured representation enables sequence alignment with video frames for zero-shot recognition.

### v3

@@ 4.2 Experimental Result
We begin with a random choice baseline and a SigLIP zero-shot baseline using mean-pooled frame embeddings. Following ViFi-CLIP Rasheed et al. (2023), each video $V_{i}$ is encoded as $\bar{\mathbf{z}}_{i}=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\mathbf{z}_{i}^{t}$, where mean-pooled embeddings $\bar{\mathbf{z}}$ is compared via cosine similarity to each name embedding $\phi_{t}(c_{j})$.
[…]
@@
Configuration Top-1 (%) $\uparrow$ Top-2 (%) $\uparrow$ Top-3 (%) $\uparrow$
SigLIP Zhai et al. (2023) (mean-pool) $22.94$ $42.20$ $63.70$
+ DTW Alignment $25.72$ $45.99$ $66.26$
+ Context Augmentation $30.07$ $52.67$ $70.49$
+ Signal Smoothing $30.29$ $53.01$ $70.27$
@@
Baseline Comparisons
To further evaluate the generalization power of SigLIP without any temporal information, we evaluate SigLIP performance by the similar context augmentation technique we used in our framework and with mean-pooling the bag of sub-actions and comparing the mean-pooled representations with video representation (See Table 3). In all of these baselines, video frame representations are mean-pooled. We further investigate classification performance under randomized and reversed sub-action orders at Table 4
Configuration Top-1 (%) $\uparrow$ Top-2 (%) $\uparrow$ Top-3 (%) $\uparrow$
SigLIP (mean-pool) $22.94$ $42.20$ $63.70$
SigLIP (mean-pool) w/ Context Augmentation $28.51$ $50.89$ $70.38$
SigLIP (mean-pool) w/ Bag-of-Words $29.06$ $51.34$ $69.49$
ActAlign (ours) $30.29$ $53.01$ $70.27$
Table 3: Comparison with baselines under context-rich prompting. Context augmentation enriches action names (as opposed to sub-actions in Table 2) with domain names, while bag-of-words mean-pools sub-action descriptions into a unified representation.
Sub-action Baselines
To evaluate the impact of sub-actions on alignment, we test DTW alignment performance under two perturbations: randomizing sub-actions and reversing their order on short-fixed sub-actions. Results are shown in Table 4.
Configuration Top-1 (%) $\uparrow$ Top-2 (%) $\uparrow$ Top-3 (%) $\uparrow$
Reversed Order $26.39$ $47.66$ $67.04$
Randomized Order (5 Trials) $26.63\pm 0.14$ $47.05\pm 0.29$ $67.00\pm 0.25$
Normal $27.06$ $46.88$ $66.82$
Table 4: Comparison with sub-action baselines. We evaluate classification performance by randomizing and reversing the order of short-fixed sub-actions due to their larger sequence size.
@@
Prompt Description Top-1 (%) $\uparrow$
Short‐fixed 2-word, fixed 10 sub-actions $27.06$
Context‐rich context-rich, variable-length 30.29
@@
Table 6 shows LLM-generated sub-action sequences for two Figure skating tactics under our context-rich prompting. The scripts highlight ordered, salient steps enabling precise temporal matching.
Sub-action Script
Biellmann Spin
1. Begins upright spin on one foot with arms extended and free leg behind, 2. Gradually pulls free leg upward behind the back using both hands, 3. Raises the free leg above head level while arching the back dramatically, 4. Grasps the blade of the free skate with both hands overhead, 5. Extends spinning leg vertically while maintaining centered spin on skating foot, 6. Maintains high-speed rotation with body in extreme vertical split position
Flying Camel Spin
1. Skater glides forward with arms extended and knees bent in preparation, 2. Performs a powerful jump off the toe pick while swinging free leg upward, 3. Rotates mid-air with body extended horizontally like a ’T’ shape, 4. Lands on one foot directly into a camel spin position with torso parallel to ice, 5. Extends free leg backward and arms outward while spinning on the skating leg, 6. Maintains fast, centered rotation in the horizontal camel position
Table 6: LLM-generated sub-action scripts for figure skating tactics. Shown for the Biellmann Spin and Flying Camel Spin examples in Figure 4, these sequences are generated using context-rich prompting and provide semantically detailed, temporally ordered steps for alignment in our zero-shot framework.
@@ 5 Conclusion
We show that contrastive image–language models establish a surprisingly strong baseline for zero-shot fine-grained video classification, even when used with simple mean-pooling. To fully leverage this capability in extremely fine-grained settings, we propose ActAlign, a novel zero-shot, domain-general, and model-agnostic framework that revisits the classic Dynamic Time Warping (DTW) algorithm to cast video classification as a sequence-alignment problem. By aligning video frames with LLM-generated sub-action scripts, ActAlign introduces temporal structure into contrastive models without requiring any video–text training or fine-tuning. Evaluated on the highly challenging and diverse ActionAtlas Salehi et al. (2024) benchmark, our method achieves state-of-the-art performance, outperforming both CLIP-style baselines and billion-parameter video–language models.
@@ B.1 System Configuration
B.2 Model Settings
We use the publicly available SigLIP model from Google (siglip-so400m-patch14-384), which supports variable input resolutions. Images are embedded using a ViT-based vision encoder with a patch size of $14\times 14$ and an output embedding dimension of $d=384$. Sub-actions are encoded using the corresponding frozen SigLIP text encoder. All embeddings are $\ell_{2}$-normalized before computing cosine similarities.
B.3 Runtime Performance
@@
B.4 Scalability Analysis
From an algorithmic standpoint, the scalability is primarily determined by the dynamic time warping (DTW) alignment between the visual frame sequence and sub-action scripts. The overall complexity scales as $\mathcal{O}(NMT)$, where $N$ is the number of candidate classes, $M$ is the average number of sub-actions per class, and $T$ is the number of video frames. In practice, $M$ remains small (on the order of tens) and independent of video length. Because the method operates entirely in a frozen embedding space and requires no gradient updates, it avoids the substantial memory and computational overhead associated with training large video-language models. As a result, ActAlign scales gracefully to longer videos and larger label vocabularies. Moreover, since alignment computations for different candidate classes and frame segments are independent, the framework is naturally amenable to parallelization across GPUs or distributed compute resources, offering a straightforward path for further acceleration in large-scale or real-time settings.
B.5 Inference Pipeline Overview
@@ Appendix C More Related Works
Image–Language Transfer for Video Tasks. Image–language models pretrained on large-scale image–text pairs exhibit strong open-set recognition capabilities. This motivates their adaptation for video tasks. Common strategies include adding temporal modules Wang et al. (2022); Ni et al. (2022); Rasheed et al. (2023); Liu et al. (2024) or introducing learnable prompts Ju et al. (2022); Lin et al. (2022). However, as noted by Rasheed et al. (2023), fine-tuning often compromises the model’s open-set generalization. To preserve zero-shot capabilities, recent work explores training-free adaptations using temporal prompting methods Ahmad et al. (2024); Phan et al. (2024); Yu et al. (2025). For example, TEAR Bosetti et al. (2025) compares mean-pooled frame embeddings with LLM-generated action descriptions for each class. While effective, these approaches either degrade generalization through fine-tuning or ignore temporal structure via mean-pooling - limiting performance on fine-grained and zero-shot video classification.
@@ Appendix D Prompt Templates and Generation Settings
This setting produces terse outputs (e.g., “tie arms,” “deliver knee”) that often lack domain-specific context. The descriptions are directly adopted from the orginal ActionAtlas Salehi et al. (2024) dataset.
D.1.1 Template
@@
This setting encourages richer descriptions with more related keywords grounded in the activity domain. Context-rich prompting strategy significantly improves textual grounding during DTW alignment.
D.2.1 Template
@@ D.3 Examples
Table 7 compares the two prompting strategies for a skateboarding trick example. The corresponding affinity matrices are shown in Figure 6.
@@ Appendix E Dataset Statistics
Table 8 summarizes key statistics of the ActionAtlas Salehi et al. (2024) dataset, which we extended to enable alignmnet-based classification. The dataset contains 898 video clips spanning 558 unique tactics across 56 sports domains. Each video is paired with 4 to 6 fine-grained candidate actions (5.25 in average per video). Videos vary significantly in length, with a mean of 173 frames (30 FPS), ranging from extremely short clips (3 frames) to multi-second sequences (up to 900 frames). Notably, ActionAtlas is the most diverse and challenging fine-grained action recognition benchmark to the best of our knowledge. It poses a major challenge in fine-grained recognition: human top-1 accuracy is just 61.64%.
@@
Figure 8: Our pipeline uses an LLM to generate structured sub-action sequences for each fine-grained candidate action in ActionAtlas Salehi et al. (2024) data sample. This structured representation enables sequence alignment with video frames for zero-shot recognition.
@@
Appendix F Performance within Domains
We evaluate our approach on ActionAtlas, which covers over 50 sports domains while including extremely fine-grained actions, to assess its generality. Specifically, we focus on the four largest domains in the dataset and compare performance against the SigLIP baseline (Table 10). Our method consistently outperforms the baseline across all domains.
Domain Samples ActAlign (Ours) SigLIP Baseline $\Delta$ ($p$)
Top-1 Top-2 Top-3 Top-1 Top-2 Top-3 Top-1 Top-2 Top-3
Soccer 168 27.38 52.98 70.83 23.81 45.83 63.10 +3.57 +7.15 +7.73
Basketball 106 36.79 54.72 70.75 17.92 41.51 58.49 +18.87 +13.21 +12.26
Cheerleading 61 22.95 55.74 67.21 22.95 37.70 60.66 +0.00 +18.04 +6.55
Wrestling 60 26.67 48.33 61.67 16.67 36.67 66.67 +10.00 +11.66 -5.00
Mean 28.45 52.94 67.62 20.34 40.43 62.23 +8.11 +12.52 +5.39
Table 10: Top-$k$ accuracy (%) comparison between ActAlign and SigLIP baseline across four largest ActionAtlas domains. ActAlign demonstrates domain-agnostic improvements across all $k$ levels. $\Delta$ denotes the absolute percentage-point gain of ActAlign over the baseline. Bold indicates the higher result for each domain and metric.
Appendix G Additional Qualitative Example
To illustrate how different phrasings of sub-actions can influence alignment behavior, we present an example of tie-down roping trick in Rodeo. Figure 9 shows the corresponding DTW alignment paths, and Table 11 lists the sub-action scripts used in each case.
(a) Video: Performing tie-down roping in Rodeo.
(b) Short-fixed sub-actions, $\hat{\gamma}=0.96$ $\checkmark$
(c) Context-rich sub-actions, $\hat{\gamma}=0.84$ $\checkmark$
Figure 9: DTW alignment paths comparison for a correctly predicted sample between short-fixed and context-rich sub-actions. The sub-action scripts are provided in Table 11.
Step Short-Fixed Script Context-Rich Script
1 Pursue calf Cowboy and horse launch from the roping box as the calf bolts across the arena.
2 Swing lasso Roper swings the rope fluidly and hurls it in a clean loop that snags the calf’s neck mid-gallop.
3 Catch neck Rope tightens as horse halts and backs to maintain tension, stopping the calf abruptly.
4 Dismount quickly Cowboy dismounts in one jump while horse holds the rope taut, sprinting to the flailing calf.
5 Reach calf Flips the calf onto its side with a quick lift-and-roll maneuver in the loose arena footing.
6 Flip calf Ties three of the calf’s legs with a piggin’ string in a swift, practiced knot.
7 Tie legs Throws hands up to signal completion and steps away as the calf remains bound for six seconds.
8 Raise hands
9 Step back
10 Stop time
Table 11: Comparison of short-fixed and context-rich sub-action scripts for tie-down roping in Rodeo. The DTW alignment paths for both instances are shown in Figure 9
Appendix H VLM Evaluation Settings
We provide here the detailed evaluation settings for the video–language models reported in Table 1 of the main paper.
Our evaluation strictly follows the official ActionAtlasSalehi et al. (2024) protocol and public leaderboard. Each model is evaluated using multiple-choice question–answer pairs, where the model must select from a fixed set of candidate actions. As part of the standardized pre-processing pipeline, all audio signals, speech transcripts, and any textual cues visible in the video are discarded or blurred prior to inference. This ensures that evaluation reflects purely visual understanding rather than reliance on auxiliary textual or audio information.
Additionally, ActionAtlas provides a comprehensive analysis of how performance varies with different frame rates. The results below correspond to the best-performing frame-rate configuration for each model as reported in the official leaderboard. The number of frames, token configuration, and Top-1 accuracy for each model are shown below:
Model Frames Tokens Top-1 Acc. (%)
Qwen2-VL-7B 16 $8\times 576$ 30.24
VideoLLaMA 16 $16\times 256$ 22.71
VideoChat2 64 $64\times 196$ 21.27
mPLUG-Owl-Video 16 $16\times 256$ 19.49
LLaVA-Next-Video-7B 64 $64\times 144$ 22.90
Table 12: Evaluation settings and results for baseline VLMs on the ActionAtlas benchmark.
This table reproduces the official ActionAtlas leaderboard conditions under which we report all baseline results in the main text.